shearrahg:

Analyze the diagrams below and complete the instructions that follow. 2004-06-01-03-00_files/i0020000.jpg Given the calculations shown, find the dimensions of the prism

Vocaloid:



Vocaloid:

lateral area = perimeter of base * height I will make a simplifying assumption that perimeter of the base is 40 and height is 7 according to the first calculation. therefore: perimeter = 2(w+h) = 40 now, using the second calculation, the base is 100, so w*h = 100 now you have two equations in two unknowns 2(w+h) = 40 and w*h = 100 so you can solve for w and h. as a hint you can solve equation 2 for either w or h, and plug into the first equation to eliminate one of the varibles